OpenStudy (anonymous):

Best Response & Fan.! Please Help.! 2z - 3y - z + y. Simplify The Expression.

OpenStudy (anonymous):

Add like terms :) 2z + (-z) = z -3y + y = -2y Simplified answer: z -2y

OpenStudy (anonymous):

Hah. I Kinda Understand. That's Somewhat What My Notes Showed. I Didn't Understand But I Know That's Right.c: @pabloecortez07

OpenStudy (anonymous):

Well, if you have a similar question, just ask, I'll try to help out

OpenStudy (anonymous):

Okay, Thanks So Much. :D @pabloecortez07

OpenStudy (anonymous):